OpenStudy (jenniferjuice):

What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)? H-E-L-P

OpenStudy (anonymous):

Using y=mx+c to make an equation. First by finding gradient,-3+3/1-(-5) = -1 Find the y-intercept(c) by using any point and substitute into y=mx+c . -3=-1(1)+c c=-4 Thus the equation will be y=-x-4 When y=0, -x-4=0 -x=4 x=-4 Hope this helps!

OpenStudy (anonymous):

Using y=mx+c to make an equation. First by finding gradient,-3+3/1-(-5) = -1 Find the y-intercept(c) by using any point and substitute into y=mx+c . -3=-1(1)+c c=-4 Thus the equation will be y=-x-4 When y=0, -x-4=0 -x=4 x=-4 Hope this helps!